# Holdout-Loss-Based Data Selection for LLM Finetuning via In-Context Learning

### v1

@@ Abstract
Fine-tuning large pretrained language models is a common approach for aligning them with human preferences, but noisy or off-target examples can dilute supervision. While small, well-chosen datasets often match the performance of much larger ones, systematic and efficient ways to identify high-value training data remain underexplored. Many current methods rely on heuristics or expensive retraining. We present a theoretically grounded, resource-efficient framework for data selection and reweighting. At its core is an In-Context Approximation (ICA) that estimates the holdout loss a model would incur after training on a candidate example by conditioning on a small, curated holdout set in context. ICA requires no reference model and no additional finetuning. Under a local linearization, ICA is equivalent to a first-order update toward the holdout optimum, motivating its use as a proxy for data value. We derive per-example weights from ICA scores, dynamically reweighting gradient updates as model parameters evolve. Across SFT, DPO, and SimPO, and over diverse backbones and datasets, ICA-based reweighting consistently improves model alignment with minimal overhead. We analyze sensitivity to score update frequency and the choice of $k$ holdout examples for in-context demonstrations, and note limitations for rapidly drifting on-policy updates, highlighting directions for future work. Code and prompts will be released.
@@ 1 Introduction
Fine-tuning has become the standard approach for adapting large pretrained language models to downstream applications and aligning them with human intent (Wei et al., 2021; Ouyang et al., 2022). This process typically starts with supervised fine-tuning (SFT) on instruction-response pairs (Wei et al., 2021), followed by preference-based alignment using pairwise preference data, e.g., RLHF (Christiano et al., 2017; Ouyang et al., 2022), DPO (Rafailov et al., 2023), and simPO (Meng et al., 2024). Since fine-tuning effectively steers the pretrained model toward desired behaviors, the quality of training data plays a central role: high-quality examples provide clear alignment signals, while noisy or inconsistent data can severely degrade performance.
However, training data for fine-tuning, typically collected from human annotators or model-generated outputs, often contain errors, inconsistencies, or redundancies. For example, Gao et al. (2024) report that 20–40% of preference pairs in LM alignment are noisy, and that alignment performance is highly sensitive to such noise. In addition, recent studies have shown that smaller but carefully curated datasets can yield alignment performance comparable to much larger ones (Zhou et al., 2023; Chen et al., 2023). Despite the recognized importance and promise of effective data selection for fine-tuning, systematic yet principled approaches remain underexplored.
@@
In this work, we propose a computationally efficient and theoretically grounded framework for data selection in model fine-tuning. Following prior work (Mindermann et al., 2022; Xia et al., 2024; Calian et al., 2025), we aim to select a subset of training data such that the model trained on it minimizes the holdout loss, i.e., the loss on this holdout set. Each example’s value is quantified by the holdout loss the model would incur if trained with that example. Computing this naively is intractable, so we build on a tractable approximation introduced by RHO-Loss (Mindermann et al., 2022), which estimates the holdout loss from a Bayesian perspective. Nonetheless, efficiently computing this approximation remains challenging, as it would normally require retraining on updated training and holdout sets at each step.
RHO-Loss mitigates per-step retraining but requires a reference model trained on the holdout set. We introduce an in-context approximation (ICA) that eliminates the need for both additional fine-tuning and a fixed reference model. Building on the insight that in-context learning performs implicit fine-tuning (Dai et al., 2023), we provide the holdout set as in-context demonstrations at each training step, simulating one step of fine-tuning. The resulting data values derived from ICA, termed ICA scores, enable dynamic evaluation of each example’s utility as the model evolves. These scores are then used to reweight gradient updates during fine-tuning, prioritizing examples that most reduce holdout loss. Experiments show that training with ICA-based reweighting consistently improves model alignment for SFT, DPO, and SimPO across diverse datasets and backbone models.
@@ 3 Preliminary
$C=\{I,s(\mathbf{x}_{1},\mathbf{y}_{1}),\dots,s(\mathbf{x}_{k},\mathbf{y}_{k})\}.$
@@ 4.1 Problem Formulation
$\bar{\mathcal{D}}^{\star}=\arg\min_{\bar{\mathcal{D}}\subset{\mathcal{D}}}{\mathcal{L}}\big(\mathcal{D}_{\mathrm{ho}};\bm{\theta}^{\star}(\bar{\mathcal{D}})\big),\quad\text{where}\quad\bm{\theta}^{\star}(\bar{\mathcal{D}})=\arg\min_{\bm{\theta}}{\mathcal{L}}(\bar{\mathcal{D}};\bm{\theta}).$ (1)
@@
Equation 1 would be prohibitively expensive to solve naively, as it requires training on every candidate subset $\bar{\mathcal{D}}\subset\mathcal{D}$ and evaluating the holdout loss. Instead of solving it directly, we recast the problem in terms of quantifying the contribution of each training example to reducing holdout loss. Concretely, we assign a score to each example that reflects this contribution. These scores can then be used either to select the example that most reduces the holdout loss, or to reweight gradient updates in the optimizer according to each example’s contribution.
In the following section, we detail how to approximate these scores without performing actual training and how to compute them efficiently using in-context learning.
@@ 4.2 Computing Approximated Holdout Loss via In-Context Learning
Approximating holdout loss via Bayesian framework
@@
Accordingly, each training example’s contribution can be measured by the holdout loss in Equation 2. To compute this holdout loss without performing actual training, we adopt an approximation derived by Mindermann et al. (2022) in the general supervised learning setting, based on probabilistic modeling. Specifically, considering the negative log-likelihood as the loss function ($\ell(\mathbf{y}\mid\mathbf{x};\bm{\theta})=-\log p(\mathbf{y}\mid\mathbf{x};\bm{\theta})$), and, applying Bayes’ rule under the conditional independence assumption, the holdout loss of the model trained with a particular example can be approximated as (see Appendix A for a reproduction of the derivation in Mindermann et al. (2022))  [^1] This Bayesian framework also extends to pairwise preference data (e.g., for DPO and simPO), where $\mathbf{y}$ represents a preference pair $(\mathbf{y}_{w},\mathbf{y}_{l})$ with $\mathbf{y}_{w}\succ\mathbf{y}_{l}$, and the loss is defined correspondingly for the chosen preference learning model (e.g., Bradley-Terry as used in DPO). :
${\mathcal{L}}(\mathcal{D}_{\mathrm{ho}};\bm{\theta}^{\star}({\mathcal{D}}_{t}\cup\{(\mathbf{x},\mathbf{y})\}))\approx{\ell}(\mathbf{y}\mid\mathbf{x};\bm{\theta}^{\star}({\mathcal{D}}_{t}\cup\mathcal{D}_{\mathrm{ho}}))-{\ell}(\mathbf{y}\mid\mathbf{x};\bm{\theta}_{t})-{\mathcal{L}}(\mathcal{D}_{\mathrm{ho}};\bm{\theta}_{t})$ (3)
@@
The optimal example can then be approximately found by maximizing this score:
@@
However, RHO-Loss uses a fixed reference model rather than re-evaluating each example’s impact on the holdout loss as the model is updated. This simplification can introduce bias in estimating each example’s contribution (Wang et al., 2024). To enable an efficient and adaptive data selection criterion, we introduce an in-context approximation that removes auxiliary training entirely by approximating $\ell(\mathbf{y}\mid\mathbf{x};\bm{\theta}^{\star}({\mathcal{D}}_{t}\cup\mathcal{D}_{\mathrm{ho}}))$ via in-context learning. This technique is described in detail below.
@@
Motivated by the finding that in-context learning can implicitly perform gradient-based model updates on the provided demonstrations (Dai et al., 2023), we avoid retraining on ${\mathcal{D}}_{t}\cup\mathcal{D}_{\mathrm{ho}}$ by using the holdout set as in-context demonstrations. Specifically, we introduce the following in-context approximation (ICA) to compute the second term in equation 4 without retraining:
@@
By conditioning on the holdout set $\mathcal{D}_{\mathrm{ho}}$, this approximation effectively simulates one step of fine-tuning on the set, providing a computationally tractable estimate of the model’s state after training on ${\mathcal{D}}_{t}\cup\mathcal{D}_{\mathrm{ho}}$.
Applying ICA to equation 4 yields a computationally efficient selection criterion, which we refer to as the in-context approximation score (ICA score):
@@
We can now use the ICA score in place of the holdout-loss score in equation 5 as the data selection criterion. This approach is not only more computationally efficient, but also enables dynamic re-evaluation of each example’s impact on the holdout loss as the model evolves.
@@ 4.3 Batch Selection via a Reweighting Strategy
Concretely, consider gradient-based training (e.g., Adam or SGD) with mini-batch updates. At each iteration $t$, a batch ${B}_{t}\subset\mathcal{D}$ of size $n_{B}$ is sampled. For each example in the batch, we compute its ICA score as defined in equation 7 and convert these scores into continuous weights in the range $[0,1]$ via max-min normalization. We adopt max–min instead of the softmax used by Wang et al. (2024); Calian et al. (2025) because it preserves the relative differences between scores and avoids the exponential amplification that can distort the contribution of low- and high-scoring examples. The weights are given by
$w(\mathbf{x}_{i},\mathbf{y}_{i};\bm{\theta}_{t})=\frac{{s}(\mathbf{x}_{i},\mathbf{y}_{i};\bm{\theta}_{t})-\min_{j\in{B}_{t}}{s}(\mathbf{x}_{j},\mathbf{y}_{j};\bm{\theta}_{t})}{\max_{j\in{B}_{t}}{s}(\mathbf{x}_{j},\mathbf{y}_{j};\bm{\theta}_{t})-\min_{j\in{B}_{t}}{s}(\mathbf{x}_{j},\mathbf{y}_{j};\bm{\theta}_{t})}.$ (8)
@@
Because the ICA score, and hence the weights, evolve throughout training, this reweighting strategy adaptively adjusts each example’s contribution to the gradient updates according to the holdout loss the model would achieve if trained on that example. Algorithm 1 outlines how the ICA score is computed and used to reweight gradient updates during fine-tuning.
@@
When computing the in-context approximation (line 6 or equation 6), including the entire holdout set as demonstrations is often infeasible due to prompt length constraints. One solution is to divide the full holdout set into multiple subsets, compute the score for each, and average the results; however, this can be computationally intensive. Instead, we select the top-$k$ holdout examples most similar to each candidate via $k$-nearest neighbor (kNN) search in an embedding space. To further improve efficiency, we periodically update the scores $R$ times during training, instead of recomputing them at every iteration. Experimental results show that even with these practical approximations, our method consistently improves alignment performance.
The complete procedure incorporating these techniques is presented as Algorithm 2 in Appendix B.4. We ablate the effects of different choices of $k$ and $R$ in Section 5.3 and analyze the computational overhead of our implementation in Section 5.4.
@@ 5 Experimental Results
We apply our method to both SFT and preference-based alignment (DPO and SimPO), comparing ICA score-based reweighting to standard training (without reweighting) and to reweighting using scores computed by baseline methods, across multiple models and datasets.
@@ 5.1 Experimental Setup
Since our goal is to select a subset of training data that minimizes holdout loss, which serves as a proxy for downstream performance, we evaluate our method by measuring how closely the model’s outputs align with the test set targets. To this end, we define the win rate as the percentage of pairwise comparisons in which a response from a model trained with our method is judged closer to the target output than that of a model trained with standard (non-reweighted) or baseline methods. Comparisons are performed by GPT-4o (2024-08-06). Due to computational constraints, we train and evaluate each model only once. The standard deviation of evaluation results is small (see Appendix B.2), so reporting a single evaluation run is sufficient.
Datasets
We consider two data selection scenarios. High-quality selection prioritizes expert-level data using a smaller curated holdout set: for SFT, we use Alpaca as training data and sample high-quality holdout examples from its curated version, Alpaca-cleaned (Taori et al., 2023); for preference optimization (DPO and SimPO), we use UltraFeedback-binarized (Cui et al., 2023), which provides preference pairs with scalar quality scores, enabling the construction of high-quality holdout and test sets. Domain-relevant selection prioritizes examples relevant to a target domain using a domain-specific holdout set: for SFT, we use Yahoo_Answers_Topic [^2] https://huggingface.co/datasets/community-datasets/yahoo_answers_topics/viewer?views%5B%5D=train, and for preference optimization, we use SHP-2 (Ethayarajh et al., 2022); both datasets contain domain labels, allowing selection of a target domain for evaluating out-of-domain alignment. Details of the dataset splits (training, test, and holdout) are provided in Appendix B.1.
@@
We evaluate our approach on multiple model families and scales, including LLaMA-3-8B-Instruct, LLaMA-3-3B-Instruct, Qwen-3-8B, and Qwen-3-4B. Models are fine-tuned using two parameter-updating paradigms: full-parameter fine-tuning and parameter-efficient LoRA (Hu et al., 2021). Due to space constraints, detailed training configurations are provided in Appendix B.3, and LoRA fine-tuning results with our method are reported in Appendix B.8.
@@
By default, we perform full-parameter training and use the ICA score to reweight training examples. We compute ICA scores using the two practical techniques described in Section 4.3 (see Appendix B.4 for additional implementation details). Specifically, for each candidate, we use the top $k=3$ holdout examples as in-context demonstrations and update scores for all training examples $R=1$ time (computing scores for all training examples only at the initialization step $t=0$). When selecting the top $k$ holdout examples in embedding space, we adopt all-mpnet-base-v2 (Reimers and Gurevych, 2019) to compute embeddings.
@@
We compare our method with standard training (without reweighting) and with reweighting using scores from the following methods in place of the ICA score: (1) RHO-Loss (Mindermann et al., 2022), which approximates the holdout loss score using a model trained on the holdout set; (2) One-shot learning (Li et al., 2023b), which scores each candidate as the difference between the one-shot loss with the candidate included as context and the zero-shot loss without it. Detailed formulas for computing scores with these two baseline methods are provided in Appendix B.5.

### v2

@@ Abstract
Fine-tuning large pretrained language models is a common approach for aligning them with human preferences, but noisy or off-target examples can dilute supervision. While small, well-chosen datasets often match the performance of much larger ones, systematic and efficient ways to identify high-value training data remain underexplored. Many current methods rely on heuristics or expensive retraining. We present a principled, resource-efficient framework for data selection and reweighting. At its core is an In-Context Approximation (ICA) that estimates the holdout loss a model would incur after training on a candidate example by conditioning on a small, curated holdout set in context. ICA requires no reference model and no additional finetuning. We define the resulting estimate as the ICA score, and derive per-example weights that dynamically reweight gradient updates as model parameters evolve. Across SFT, DPO, and SimPO, and over diverse backbones and datasets, ICA-based reweighting consistently improves model alignment with minimal overhead. We analyze sensitivity to score update frequency and the number of in-context holdout examples. We also discuss limitations in rapidly drifting on-policy settings, highlighting directions for future work. Code and prompts will be released.
@@ 1 Introduction
Fine-tuning has become the standard approach for adapting large pretrained models to downstream applications and aligning them with human intent (Wei et al., 2021; Ouyang et al., 2022). This process typically starts with supervised fine-tuning (SFT) on instruction-response pairs (Wei et al., 2021), followed by preference-based alignment using pairwise preference data, e.g., RLHF (Christiano et al., 2017; Ouyang et al., 2022), DPO (Rafailov et al., 2023), and simPO (Meng et al., 2024). Since fine-tuning effectively steers the pretrained model toward desired behaviors, the quality of training data plays a central role: high-quality examples provide clear alignment signals, while noisy or inconsistent data can severely degrade performance.
However, training data for fine-tuning, typically collected from human annotators or model-generated outputs, often contain errors, inconsistencies, or redundancies. For example, Gao et al. (2024) report that 20–40% of preference pairs in LM alignment are noisy, and that alignment performance is highly sensitive to such noise. In addition, recent studies have shown that smaller but carefully curated datasets can yield alignment performance comparable to much larger ones (Zhou et al., 2023; Chen et al., 2023). Despite the recognized importance and promise of effective data selection for fine-tuning, computationally efficient and principled approaches remain underexplored.
@@
In this work, we propose a principled, resource-efficient framework for data selection in model fine-tuning. Following prior work (Mindermann et al., 2022; Xia et al., 2024; Calian et al., 2025), we aim to select a subset of training data such that the model trained on it minimizes the holdout loss, i.e., the loss on the holdout set. Each example’s value is quantified based on the holdout loss the model would incur if trained with that example. Computing this naively is intractable, so we build on a framework introduced by RHO-Loss (Mindermann et al., 2022), which uses Bayesian probability theory to derive a more tractable expression for the holdout loss. Nonetheless, efficiently computing this expression remains challenging, as evaluating it would normally require retraining at each step.
RHO-Loss mitigates per-step retraining by relying on a separate reference model trained on the holdout set. We introduce an in-context approximation (ICA) that eliminates the need for both additional fine-tuning and a fixed reference model. Motivated by the insight from (Dai et al., 2023), we provide the holdout set as in-context demonstrations at each training step to approximate the effects of explicit parameter updates on model performance. The resulting data values derived from ICA, termed ICA scores, enable dynamic evaluation of each example’s utility as the model evolves. These scores are then used to reweight gradient updates during fine-tuning, prioritizing examples that most reduce holdout loss. Experiments show that training with ICA-based reweighting consistently improves model alignment for SFT, DPO, and SimPO across diverse datasets and backbone models.
@@ 3 Preliminary
$C=\{I,s(\mathbf{x}_{1},\mathbf{y}_{1}),\dots,s(\mathbf{x}_{k},\mathbf{y}_{k})\}$
where $s(\cdot,\cdot)$ is a formatting function (Dong et al., 2022).
@@ 4.1 Problem Formulation
$\begin{split}\bar{\mathcal{D}}^{\star}&=\arg\min_{\bar{\mathcal{D}}\subset{\mathcal{D}}}{\mathcal{L}}\big(\mathcal{D}_{\mathrm{ho}};\bm{\theta}^{\star}(\bar{\mathcal{D}})\big),\\ \quad\text{where}&\quad\bm{\theta}^{\star}(\bar{\mathcal{D}})=\arg\min_{\bm{\theta}}{\mathcal{L}}(\bar{\mathcal{D}};\bm{\theta}).\end{split}$ (1)
@@
Problem equation 1 would be prohibitively expensive to solve naively, as it requires training on every candidate subset $\bar{\mathcal{D}}\subset\mathcal{D}$ and evaluating the holdout loss. Instead of solving it directly, we recast the problem in terms of quantifying the contribution of each training example to reducing holdout loss. Concretely, we assign a score to each example that reflects this contribution. These scores can then be used either to select the example that most reduces the holdout loss, or to reweight gradient updates in the optimizer according to each example’s contribution.
In the following section, we describe an in-context learning–based approach for approximating these scores, removing the need for auxiliary retraining.
@@ 4.2 Computing Approximated Holdout Loss via In-Context Learning
Bayesian formulation of the holdout loss
@@
Accordingly, each training example’s contribution can be measured by this holdout loss.
To avoid explicit retraining when evaluating the holdout loss in equation 2, we build on the Bayesian formulation introduced by Mindermann et al. (2022). Specifically, considering the negative log-likelihood as the loss function ($\ell(\mathbf{y}\mid\mathbf{x};\bm{\theta})=-\log p(\mathbf{y}\mid\mathbf{x};\bm{\theta})$), and, applying Bayes’ rule under the conditional independence assumption, the holdout loss of the model trained with a particular example can be expressed as (see Appendix A for a reproduction of the derivation in Mindermann et al. (2022))  [^1] This Bayesian framework also extends to pairwise preference data (e.g., for DPO and simPO), where $\mathbf{y}$ represents a preference pair $(\mathbf{y}_{w},\mathbf{y}_{l})$ with $\mathbf{y}_{w}\succ\mathbf{y}_{l}$, and the loss is defined correspondingly for the chosen preference learning model (e.g., Bradley-Terry as used in DPO). :
$\begin{split}{\mathcal{L}}\big(\mathcal{D}_{\mathrm{ho}};\bm{\theta}^{\star}({\mathcal{D}}_{t}\cup\{(\mathbf{x},\mathbf{y})\})\big)={\ell}\big(\mathbf{y}\mid\mathbf{x};\bm{\theta}^{\star}({\mathcal{D}}_{t}\cup\mathcal{D}_{\mathrm{ho}})\big)-{\ell}(\mathbf{y}\mid\mathbf{x};\bm{\theta}_{t})-{\mathcal{L}}(\mathcal{D}_{\mathrm{ho}};\bm{\theta}_{t})\end{split}$ (3)
@@
The optimal example can then be found by maximizing this score:
@@
Nonetheless, RHO-Loss still incurs the cost of training a separate reference model. Moreover, using a fixed model in place of one that would be updated at each selection step can introduce bias in estimating each example’s contribution (Wang et al., 2024). To enable an efficient and adaptive data selection criterion, we introduce an in-context approximation that removes auxiliary training entirely by approximating $\ell(\mathbf{y}\mid\mathbf{x};\bm{\theta}^{\star}({\mathcal{D}}_{t}\cup\mathcal{D}_{\mathrm{ho}}))$ via in-context learning. This technique is described in detail below.
@@
Motivated by the finding that in-context learning can induce model behaviors similar to those obtained through gradient-based updates (Dai et al., 2023), we provide the holdout set as in-context demonstrations to a model trained on $\mathcal{D}_{t}$ to approximate the effect of training on the combined set ${\mathcal{D}}_{t}\cup\mathcal{D}_{\mathrm{ho}}$. Specifically, the second term in equation 4 is approximated as
@@
We call the approximation in equation 6 the in-context approximation (ICA).
Applying this approximation to equation 4 yields the in-context approximation score (ICA score):
@@
We can now use the ICA score to approximately identify the data point that maximizes the holdout-loss score in equation 5. This approach is not only more computationally efficient, but also enables dynamic re-evaluation of each example’s impact on the holdout loss as the model evolves.
@@ 4.3 Batch Selection via a Reweighting Strategy
Concretely, consider gradient-based training (e.g., Adam or SGD) with mini-batch updates. At each iteration $t$, a batch ${B}_{t}\subset\mathcal{D}$ of size $n_{B}$ is sampled. For each example in the batch, we compute its ICA score as defined in equation 7 and convert these scores into continuous weights in the range $[0,1]$ via max-min normalization. We adopt max–min instead of the softmax used by Wang et al. (2024); Calian et al. (2025) because it preserves the relative differences between scores and avoids the exponential amplification that can distort the contribution of low- and high-scoring examples. The per-example weights are given by
$\begin{split}w(\mathbf{x}_{i},\mathbf{y}_{i};\bm{\theta}_{t})=\frac{{s}(\mathbf{x}_{i},\mathbf{y}_{i};\bm{\theta}_{t})-\min_{j\in{B}_{t}}{s}(\mathbf{x}_{j},\mathbf{y}_{j};\bm{\theta}_{t})}{\max_{j\in{B}_{t}}{s}(\mathbf{x}_{j},\mathbf{y}_{j};\bm{\theta}_{t})-\min_{j\in{B}_{t}}{s}(\mathbf{x}_{j},\mathbf{y}_{j};\bm{\theta}_{t})}.\end{split}$ (8)
@@
Because the ICA score, and hence the weights, evolve throughout training, this reweighting strategy adaptively adjusts each example’s contribution to the gradient updates according to the holdout loss the model would incur if trained on that example. Algorithm 1 outlines how the ICA score is computed and used to reweight gradient updates during fine-tuning.
@@
When computing the in-context approximation in equation 6, including the entire holdout set as demonstrations is often infeasible due to prompt length constraints. A straightforward approach is to partition the holdout set into multiple subsets, compute scores for each subset, and average the results; however, this is computationally expensive. Instead, we select the top-$k$ holdout examples most similar to each candidate using $k$-nearest neighbor (kNN) search in an embedding space. In addition, rather than computing scores at every training iteration, we update them only $R$ times periodically over the course of training. Experimental results show that, despite these practical approximations, our method consistently improves alignment performance.
The complete procedure incorporating these techniques is presented as Algorithm 2 in Appendix B.4. We ablate the effects of different choices of $k$ and $R$ in Section 5.3 and analyze the computational overhead of our implementation in Section $5.$
@@ 5 Experimental Results
We apply our method to both SFT and preference-based alignment (DPO and SimPO), comparing ICA score-based reweighting to standard training (i.e., without reweighting) and to reweighting using scores computed by baseline methods, across multiple models and datasets.
@@ 5.1 Experimental Setup
Datasets
We consider two data selection scenarios. High-quality selection prioritizes expert-level data using a smaller curated holdout set: for SFT, we use Alpaca as training data and sample high-quality holdout examples from its curated version, Alpaca-cleaned (Taori et al., 2023); for preference optimization (DPO and SimPO), we use UltraFeedback-binarized (Cui et al., 2023), which provides preference pairs with scalar quality scores, enabling the construction of high-quality holdout and test sets. Domain-relevant selection prioritizes examples relevant to a target domain using a domain-specific holdout set: for SFT, we use Yahoo_Answers_Topics [^2] https://huggingface.co/datasets/community-datasets/yahoo_answers_topics/viewer?views%5B%5D=train, and for preference optimization, we use SHP-2 (Ethayarajh et al., 2022); both datasets contain domain labels, allowing selection of a target domain for evaluating out-of-domain alignment. Dataset split details are provided in Appendix B.1.
Within the high-quality data selection scenario, we additionally evaluate our method on a reasoning-intensive benchmark, GSM8K (Cobbe et al., 2021), which consists of high-quality grade-school math word problems. To test the method’s ability to identify high-quality examples in noisy data, a random fraction of chain-of-thought (CoT) examples in the training split are intentionally perturbed. Experimental setup details are provided in Section 5.4, with full results deferred to Appendix C.4.
@@
The objective of data selection in this paper is to choose a subset of training data such that a model trained on it minimizes the holdout loss. Accordingly, we evaluate our method by measuring how closely the model’s outputs align with the test set targets. To this end, we define the win rate as the percentage of pairwise comparisons in which a response from a model trained with our method is judged more semantically aligned with the target output than that of a model trained with standard (non-reweighted) or baseline methods. Comparisons are performed by GPT-4o (2024-08-06).
In the main experiments, each model is trained and evaluated once due to computational constraints. An analysis of training and evaluation variability is provided in Appendix B.2.
@@
We evaluate our approach on multiple model families and scales, including LLaMA-3-8B-Instruct, LLaMA-3-3B-Instruct, Qwen-3-8B, and Qwen-3-4B. Models are fine-tuned using two parameter-updating paradigms: full-parameter fine-tuning and parameter-efficient LoRA (Hu et al., 2021). Detailed training configurations are provided in Appendix B.3, and LoRA fine-tuning results with our method are reported in Appendix C.1.
@@
By default, we perform full-parameter training and use the ICA score to reweight training examples. We compute ICA scores using the two practical techniques described in Section 4.3 (see Appendix B.4 for additional implementation details). Specifically, for each candidate, we use the top $k=3$ holdout examples as in-context demonstrations and update scores for all training examples $R=1$ time (i.e., computing scores for all training examples only at the initialization step $t=0$). When selecting the top $k$ holdout examples in embedding space, we adopt all-mpnet-base-v2 (Reimers and Gurevych, 2019) to compute embeddings.
@@
In the main experiments, we compare our method with standard training (without reweighting) and two closely related baselines using alternative scores: (1) RHO-Loss (Mindermann et al., 2022), which approximates the holdout loss score using a model trained on the holdout set; (2) One-shot learning (Li et al., 2023b), which scores each candidate as the difference between the one-shot loss with the candidate included as context and the zero-shot loss without it. Detailed formulas for computing scores with these two methods are provided in Appendix B.5.
In the SFT setting, we further compare against LESS (Xia et al., 2024) and GREATS (Wang et al., 2024), two data selection methods based on influence functions. For a fair comparison, we adopt the original selection scheme of each method: when comparing to LESS, we select the top 5% of examples based on precomputed scores before training; when comparing to GREATS, we select the top 50% of examples within each batch for gradient updates. Comparison results are summarized in Section 5.2 and provided in full in Appendix C.3.